OpenStudy (anonymous):

Could someone please show me how to solve this? x times x^ (-1/4)

OpenStudy (jdoe0001):

\(\large \bf recall\implies x^{\color{red}{ n}}\cdot x^{\color{blue}{ m}}\implies x^{{\color{red}{ n}}+{\color{blue}{ m}}}\)

OpenStudy (anonymous):

So would the answer just stay as x^ (-1/4)? I'm just wondering what happens to the first x

OpenStudy (anonymous):

actually no its 1/12

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

Could you show me how its done?

OpenStudy (anonymous):

well i am not a good ex-plainer i did in my head

OpenStudy (mertsj):

\[x^1(x ^{\frac{-1}{4}})=x ^{1+(\frac{-1}{4})}=x ^{\frac{3}{4}}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

&thanks Vanoss for your attempt

OpenStudy (mertsj):

yw